OpenStudy (anonymous):

What is the center and radius of the circle with the given equation? x^2 + y^2 +10x - 6y = -9 Someone please help me or teach me or explain

OpenStudy (campbell_st):

do you know how to complete the square...?

OpenStudy (campbell_st):

if you group things \[(x^2 + 10x ) + (y^2 -6y) = -9\] each set of brackets on the left hand side need to contain perfect squares... and the process of completing the square is essential to solving this problem.

OpenStudy (anonymous):

No im not really good at this.

OpenStudy (campbell_st):

ok... so its easy.... the middle term is double the product of x and a number so find the number looking at the x part double the product is 10x so halve it 5x so 5 and x are the factors... so add 5^2 to get a perfect square.... this needs to be done to both sides of the equation.. so you get \[(x^2 + 10x + 25) + (y^2 - 6y ) = -9 + 25\] see if you can work on the y-part

OpenStudy (anonymous):

Oh thanks so is y=3 ?

OpenStudy (anonymous):

and would the radius be 25?

OpenStudy (campbell_st):

well 2y*? = -6y divide both sides by 2y ? = -3 so square it and add it to both sides of the equation.

OpenStudy (anonymous):

Ohhhhhhh okay -6/2 = -3y

OpenStudy (anonymous):

How can i find the radius?

OpenStudy (campbell_st):

ok... so what was the squared value you added to both sides of the equation... ?

OpenStudy (campbell_st):

ok... so the equation becomes \[(x^2 - 10x + 25) + (y^2 - 6y + 9) = -9 + 25 + 9\] factor the 2 quadratic equations.... they are perfect squares... and simplify the right hand side then the standard for of the circle is \[(x - h)^2 + (y - k)^2 = r^2\] the centre is (h, k) and the radius is r hope this helps